OpenStudy (falco276):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! segment PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (falco276):

@Gin_Ichimaru

OpenStudy (anonymous):

use this formula \[(\frac{ x + 25}{2}, \frac{ y+ (-5) }{ 2 } =(18,1)\]

OpenStudy (falco276):

????????????????

OpenStudy (anonymous):

Yeah falco Or you could actually graph it out on graph paper!

OpenStudy (falco276):

Can't. :( I was BORN to see answers, not seek out advice. lol. No wait. I'm serious. (sigh) Am confused. :(

OpenStudy (anonymous):

x = 11 y= 7 so ur answer would be (11,7)

OpenStudy (falco276):

Did you use the formula? But how would I know about the coordinates?

OpenStudy (anonymous):

ok so all you have to do is break it down. start off with the x \[\frac{ x+25 }{ 2 } =18\]

OpenStudy (falco276):

okay.

OpenStudy (anonymous):

then do the same for the y

OpenStudy (falco276):

Ok, I think understand now. :)